# From Sublinear to Linear: Fast Convergence in Deep Networks via Locally Polyak-Lojasiewicz Regions

### v1

@@ Abstract
The convergence of gradient descent (GD) on the non-convex loss landscapes of deep neural networks (DNNs) presents a fundamental theoretical challenge. While recent work has established that GD converges to a stationary point at a sublinear rate within locally quasi-convex regions (LQCRs), this fails to explain the exponential convergence rates consistently observed in practice. In this paper, we resolve this discrepancy by proving that under a mild assumption on Neural Tangent Kernel (NTK) stability, these same regions satisfy a local Polyak-Lojasiewicz (PL) condition. We introduce the concept of a Locally Polyak-Lojasiewicz Region (LPLR), where the squared gradient norm lower-bounds the suboptimality gap, prove that properly initialized finite-width networks admit such regions around initialization, and establish that GD achieves linear convergence within an LPLR, providing the first finite-width guarantee that matches empirically observed rates. We validate our theory across diverse settings, from controlled experiments on fully-connected networks to modern ResNet architectures trained with stochastic methods, demonstrating that LPLR structure emerges robustly in practical deep learning scenarios. By rigorously connecting local landscape geometry to fast optimization through the NTK framework, our work provides a definitive theoretical explanation for the remarkable efficiency of gradient-based optimization in deep learning.
Keywords: Deep learning theory, non-convex optimization, linear convergence, Polyak-Lojasiewicz condition, neural tangent kernel
MSC 2020 Subject Classification: 68T07, 90C26, 65K10
1.  Introduction
@@
The empirical success of deep learning is largely driven by a remarkable phenomenon: the surprising effectiveness of simple, gradient-based optimizers. The loss functions of deep neural networks (DNNs) are highly non-convex, and classical optimization theory suggests that first-order methods like gradient descent (GD) should be thwarted by a landscape riddled with spurious local minima and saddle points. This stark discrepancy between theory and practice has been a central puzzle. Early work began to resolve this by showing that the puzzle was more nuanced; for instance, by modeling the loss with a spin-glass Hamiltonian, choromanska2015loss showed that poor local minima are exponentially rare in high dimensions, though saddle points are plentiful. Subsequent work confirmed that for deep linear networks, all local minima are in fact global kawaguchi2016deep .
Recent advances have begun to bridge this theory-practice gap through more sophisticated analyses of local geometry and convergence conditions. an2024convergence established convergence of SGD under local Lojasiewicz conditions for deep networks, while recent work has derived local PL conditions for overparameterized linear models under relaxed assumptions xu2025local . Simultaneously, the development of specialized architectures has shown that PL-like properties can be built into network design: wang2024monotone introduced PLNet architectures that inherently satisfy the Polyak-Łojasiewicz condition. However, these approaches either require restrictive assumptions about network architecture or apply only to simplified settings like linear networks. Our work addresses this limitation by demonstrating that PL-like properties emerge naturally from the geometric structure of locally quasi-convex regions (LQCRs) in practical deep networks, without requiring architectural modifications or restrictive assumptions about infinite width.
@@
This has motivated several lines of research to explain why GD still succeeds. One line focuses on landscape characterization, arguing that for sufficiently wide networks, the geometry of the loss surface is benign. It has been shown that if a single hidden layer is wider than the dataset, almost all local minima are global nguyen2017loss , and that even when suboptimal minima exist, they often lie on connected manifolds that are non-attracting, allowing GD to escape petzka2021nonattracting . Another powerful line of research focuses on over-parameterization, proving that for sufficiently wide networks, GD with random initialization finds a global minimum, often at a linear rate du2019width ; du2018global . A key insight here is that over-parameterization ensures the associated Gram matrix remains stable throughout training, guaranteeing convergence du2018global . Importantly, these guarantees do not always require extreme network widths; some theories show that widths scaling just linearly with dataset size can be sufficient for finding global, generalizable solutions kawaguchi2020practical . Recent work by Aich2025 introduced Locally Quasi-Convex Regions (LQCRs), demonstrating that even practical, finite-width networks possess well-behaved regions around initialization where GD is guaranteed to converge to a stationary point, albeit with a sublinear rate.
@@
A crucial gap remains: a sublinear rate does not capture the fast, exponential convergence often observed in practice. This paper resolves this inconsistency. We build upon the LQCR framework to prove that GD can, and does, achieve a linear convergence rate under practical conditions. Our key insight is to connect the geometry of LQCRs to the Polyak-Lojasiewicz (PL) condition Polyak1963 ; Karimi2016 . While the PL condition is too strong to hold globally for DNNs, we show that it is sufficient for it to hold locally within the same region where optimization takes place.
@@
We formalize this by introducing the Locally Polyak-Lojasiewicz Region (LPLR). We prove that under a mild assumption on the stability of the NTK within an LQCR, the region is guaranteed to be an LPLR. Our main contributions are as follows:
  • Natural Emergence of PL Properties: Unlike recent work that designs specialized architectures to satisfy PL conditions wang2024monotone or assumes them for specific models an2024convergence , we prove that such properties emerge naturally from geometric structure in standard networks.
  • Finite-Width Guarantees: While most NTK-based analyses require infinite or exponentially large widths, our local stability condition holds for moderately overparameterized networks, making our results practically relevant.
  • General Architecture Applicability: Our framework applies to standard deep networks without architectural modifications, unlike specialized approaches wang2024monotone that require specific network designs.
  • Bridge Theory-Practice Gap: We provide the first finite-width explanation for the linear convergence rates consistently observed in practice, resolving a fundamental discrepancy in deep learning theory.
This work provides a more complete picture of DNN optimization, showing not only that GD converges, but that it converges fast, and provides a rigorous theoretical basis for this widely observed phenomenon. The paper is structured as follows. Section 2 reviews related work. Section 4 formalizes our framework. Our main theoretical results are presented in Section 5. We then empirically validate these findings in Section 6, and conclude in Section 7.
2.  Related Work
@@ 2.1 Local Convergence Analysis and PL-type Conditions
The classical Polyak-Lojasiewicz condition has experienced a renaissance in deep learning theory. an2024convergence ; Polyak1963 ; delarue2025genericity proved convergence of stochastic gradient descent under a local Lojasiewicz condition, showing that even approximate versions of the PL property can drive convergence in non-convex settings. This builds on the foundational work of Karimi2016 , who demonstrated that the PL condition provides a unifying framework for understanding linear convergence across diverse optimization algorithms Karimi2016 ; Necoara2015 .
Recent Architectural Developments: wang2024monotone introduced a fundamentally different approach by designing networks (PLNet) that inherently satisfy the Polyak-Łojasiewicz condition by construction. While innovative, this approach requires specialized architectures and cannot explain why standard networks exhibit fast convergence.
Linear Network Analysis: Recent theoretical progress has established local PL conditions for two-layer linear networks under general losses and relaxed assumptions on step size, width, and initialization xu2025local . Moreover, mild over-parameterization alone suffices for linear convergence in two-layer nets, as shown by zhou2021local . However, this analysis is limited to linear networks and does not extend to the non-linear case that dominates practical applications.
@@ 2.2 Neural Tangent Kernel Theory and Finite-Width Analysis
The Neural Tangent Kernel framework, introduced by Jacot2018 , revolutionized the theoretical understanding of neural network training by showing that infinitely wide networks behave as kernel methods. This insight has been crucial for proving global convergence guarantees, but typically requires unrealistic infinite-width assumptions Jacot2018 ; novak2022fast .
Finite-Width Extensions: Recent work has sought to extend NTK insights to finite-width settings. AllenZhu2019 showed that sufficiently wide networks can maintain NTK stability, while Arora2019a characterized when finite networks approximate the infinite-width behavior Arora2019a ; AllenZhu2019 ; frei2022proxy . However, these results often require exponentially large widths to provide meaningful guarantees.
@@ 2.3 Landscape Geometry and Optimization Dynamics
Understanding the geometric structure of neural network loss landscapes has been crucial for explaining optimization success. choromanska2015loss showed that poor local minima are exponentially rare using spin-glass models, while kawaguchi2016deep proved that all local minima are global for deep linear networks.
Local Geometric Analysis: Recent work has developed general mathematical frameworks for loss landscapes in over-parameterized systems, providing modern views of efficient optimization. The LQCR framework Aich2025 contributes to this line by providing concrete geometric conditions that ensure both convergence and fast rates.
@@
3.  Notation
In this section, we introduce the notation used throughout our analysis. We consider an optimization problem for deep neural networks with the following symbols: $f_{\theta}$, $\theta$, $p$, $D$, $m$, $x$, $d$, $y$, $n$, $\mathcal{L}$, $\ell$, $\mathcal{R}$, $\eta$, $t$, $L$, $L_{\Theta}$, $\alpha$, $\gamma$, $\mu$, $\mathcal{L}^{*}$, $\mathcal{L}_{\mathcal{R}}^{*}$, $J_{\theta}$, $\mathbf{\Theta}_{\theta}$, and $\lambda_{\min}$.
@@ 3.  Notation
4.  From Local Geometry to the Polyak-Lojasiewicz Condition
The PL condition, first introduced by Polyak Polyak1963 , is a powerful tool for proving linear convergence on non-convex functions. However, the global PL condition, defined as
$\frac{1}{2}||\nabla\mathcal{L}(\theta)||^{2}\geq\mu(\mathcal{L}(\theta)-\mathcal{L}^{*})$ (1)
@@ 4.  From Local Geometry to the Polyak-Lojasiewicz Condition
Definition 4.1 (Locally Polyak-Lojasiewicz (LPL) Region)
@@
$\frac{1}{2}||\nabla\mathcal{L}(\theta)||^{2}\geq\mu(\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*})\quad\forall\theta\in\mathcal{R}$ (2)
@@
The existence of such an LPLR is not arbitrary; it is a direct consequence of the underlying geometry of the region $\mathcal{R}$. Specifically, this analysis builds upon the concept of a Locally Quasi-Convex Region (LQCR) introduced in Aich2025 . An LQCR is a region with bounded negative curvature and a guaranteed descent direction, properties that ensure iterates make steady progress.
Definition 4.2 (Locally Quasi-Convex Region Aich2025 )
@@
  1. i.
    Curvature Bound: $\mathcal{L}(\phi)\geq\mathcal{L}(\theta)+\nabla\mathcal{L}(\theta)^{\top}(\phi-\theta)-\frac{\alpha}{2}||\phi-\theta||^{2}$.
  2. ii.
    Descent Condition: If $\mathcal{L}(\phi)<\mathcal{L}(\theta)$, then $\langle\nabla\mathcal{L}(\theta),\phi-\theta\rangle\geq\gamma||\phi-\theta||||\nabla\mathcal{L}(\theta)||$.
The critical link between the geometry of an LQCR and the functional property of an LPLR is the Neural Tangent Kernel (NTK) Jacot2018 . The empirical NTK, $\Theta_{\theta}$, is defined from the network Jacobian $J_{\theta}(x)$ as $\Theta_{\theta}(x,x^{\prime})=J_{\theta}(x)J_{\theta}(x^{\prime})^{\top}$. For our analysis to hold, we assume that this kernel remains well-behaved within the LQCR. This local stability is expected in many modern deep learning scenarios, including sufficiently overparameterized or wide networks, and when using standard initialization schemes and constrained learning rates Arora2019a ; AllenZhu2019 ; Aich2025 .
@@
$\mathcal{L}(\theta)-\mathcal{L}_{\mathcal{R}}^{*}\leq\frac{1}{2\lambda_{\min}(\Theta_{\theta})}||\nabla\mathcal{L}(\theta)||^{2}$ (3)
@@
5.  Main Theoretical Results

### v2

@@ Abstract
Gradient descent (GD) on deep neural network loss landscapes is non-convex, yet often converges far faster in practice than classical guarantees suggest. Prior work shows that within locally quasi-convex regions (LQCRs), GD converges to stationary points at sublinear rates, leaving the commonly observed near-exponential training dynamics unexplained. We show that, under a mild local Neural Tangent Kernel (NTK) stability assumption, the loss satisfies a PL-type error bound within these regions, yielding a Locally Polyak–Łojasiewicz Region (LPLR) in which the squared gradient norm controls the suboptimality gap. For properly initialized finite-width networks, we show that under local NTK stability this PL-type mechanism holds around initialization and establish linear convergence of GD as long as the iterates remain within the resulting LPLR. Empirically, we observe PL-like scaling and linear-rate loss decay in controlled full-batch training and in a ResNet-style CNN trained with mini-batch SGD on a CIFAR-10 subset, indicating that LPLR signatures can persist under modern architectures and stochastic optimization. Overall, the results connect local geometric structure, local NTK stability, and fast optimization rates in a finite-width setting.
Machine Learning, ICML
1 Introduction
@@
The empirical success of deep learning is largely driven by a remarkable phenomenon: the surprising effectiveness of simple, gradient-based optimizers. The loss functions of deep neural networks (DNNs) are highly non-convex, and classical optimization theory suggests that first-order methods like gradient descent (GD) should be thwarted by a landscape riddled with spurious local minima and saddle points. This stark discrepancy between theory and practice has been a central puzzle. Early work began to resolve this by showing that the puzzle was more nuanced; for instance, by modeling the loss with a spin-glass Hamiltonian, (Choromanska et al., 2015) showed that poor local minima are exponentially rare in high dimensions, though saddle points are plentiful. Subsequent work confirmed that for deep linear networks, all local minima are in fact global (Kawaguchi, 2016).
Recent advances have continued to narrow the theory-practice gap through refined local analyses of optimization dynamics. (An and Lu, 2024) established convergence guarantees for SGD under local Łojasiewicz-type conditions, and related progress has derived local PL-style properties for overparameterized linear models under relaxed assumptions (Xu et al., 2025). In parallel, architectural approaches have shown that PL-like behavior can be achieved by design: (Wang et al., 2024) introduced PLNet architectures that satisfy a Polyak-Łojasiewicz condition by construction. While these lines of work provide valuable guarantees, they typically either impose conditions tailored to specific architectures or focus on simplified settings (e.g., linear networks). Our goal is different: we aim to explain how PL-like mechanisms can arise within standard deep networks from the same local geometric structure that already underlies known convergence guarantees.
@@
This objective aligns with several influential strands of theory explaining why GD succeeds despite non-convexity. One line focuses on landscape characterization, arguing that for sufficiently wide networks the geometry is benign. For example, if a single hidden layer is wider than the dataset, almost all local minima are global (Nguyen and Hein, 2017), and even when suboptimal minima exist, they can lie on connected manifolds that are non-attracting, allowing GD to escape (Petzka and Sminchisescu, 2021). Another powerful line focuses on over-parameterization, proving that sufficiently wide networks with random initialization can reach global minima, often with linear-rate guarantees (Du and Hu, 2019; Du et al., 2018). A key insight in these results is the stability of an associated Gram matrix throughout training, which controls the dynamics and enables sharp rates (Du et al., 2018). Importantly, these guarantees do not always require extreme widths; some theories suggest that widths scaling only linearly with dataset size can suffice for finding global, generalizable solutions (Kawaguchi and Huang, 2020). Complementing these global-style analyses, (Aich et al., 2025) introduced Locally Quasi-Convex Regions (LQCRs), showing that practical, finite-width networks possess well-behaved regions around initialization in which GD is guaranteed to converge to a stationary point, albeit with a sublinear rate.
@@
A crucial gap remains: sublinear rates do not capture the fast, near-exponential decay of training objectives frequently observed in practice. This paper addresses that gap by strengthening the LQCR picture. We show that, under a mild local condition, GD can achieve a linear convergence rate within the same region where the iterates are provably confined. Our key insight is to connect the geometry of LQCRs to the Polyak-Łojasiewicz (PL) condition (Polyak, 1963; Karimi et al., 2016). While a global PL inequality is too strong to expect for general DNN objectives, it is enough for it to hold locally along the optimization trajectory, i.e., in the region where the algorithm actually operates.
@@
We formalize this by introducing the Locally Polyak-Łojasiewicz Region (LPLR). We prove that under a mild assumption on the stability of the Neural Tangent Kernel (NTK) (Jacot et al., 2018) within an LQCR, the loss satisfies a PL-type error bound within the region, yielding linear convergence of GD inside that region, as long as the iterates remain within it. Our main contributions are as follows:
  1. 1.
    Natural emergence of PL-type behavior: In contrast to approaches that build PL structure into specialized architectures (Wang et al., 2024) or assume PL/Lojasiewicz conditions directly (An and Lu, 2024), we derive a local PL-type guarantee from the intrinsic geometric structure of LQCRs in standard deep networks.
  2. 2.
    Finite-width rates under a local stability condition: Rather than requiring global NTK stability or asymptotic infinite-width limits, we identify a local NTK stability assumption sufficient to induce an LPLR and yield linear-rate convergence within it, keeping the result aligned with practical finite-width training.
  3. 3.
    Architecture-agnostic framework: The LPLR mechanism applies to standard deep networks without architectural modifications, complementing design-based approaches such as PLNet (Wang et al., 2024).
  4. 4.
    Sharper explanation of fast optimization within provable regions: Building on the LQCR guarantee of convergence to a stationary point (Aich et al., 2025), our results upgrade the predicted rate inside these regions from sublinear to linear under local NTK stability, bringing theory closer to the rapid optimization dynamics commonly observed.
Overall, this work provides a more complete picture of DNN optimization: not only why GD converges in well-behaved regions around initialization, but also when and why it converges rapidly. The remainder of the paper is organized as follows. Section 2 reviews related work. Section 3 introduces notation and preliminaries. Section 4 formalizes our framework. Section 5 presents the main theoretical results. Section 6 provides empirical validation, and finally, Section 7 concludes with directions for future work.
2 Related Work
@@ 2.1 Local Convergence Analysis and PL-type Conditions
The classical Polyak-Lojasiewicz condition has experienced a renaissance in deep learning theory. (An and Lu, 2024; Polyak, 1963; Delarue, 2025) proved convergence of stochastic gradient descent under a local Lojasiewicz condition, showing that even approximate versions of the PL property can drive convergence in non-convex settings. This builds on the foundational work of (Karimi et al., 2016), who demonstrated that the PL condition provides a unifying framework for understanding linear convergence across diverse optimization algorithms (Karimi et al., 2016; Necoara et al., 2019).
Recent Architectural Developments: (Wang et al., 2024) introduced a fundamentally different approach by designing networks (PLNet) that inherently satisfy the Polyak-Łojasiewicz condition by construction. While innovative, this approach requires specialized architectures and cannot explain why standard networks exhibit fast convergence.
Linear Network Analysis: Recent theoretical progress has established local PL conditions for two-layer linear networks under general losses and relaxed assumptions on step size, width, and initialization (Xu et al., 2025). Moreover, mild over-parameterization alone suffices for linear convergence in two-layer nets, as shown by (Zhou et al., 2021). However, this analysis is limited to linear networks and does not extend to the non-linear case that dominates practical applications.
@@ 2.2 Neural Tangent Kernel Theory and Finite-Width Analysis
The Neural Tangent Kernel framework, introduced by (Jacot et al., 2018), revolutionized the theoretical understanding of neural network training by showing that infinitely wide networks behave as kernel methods. This insight has been crucial for proving global convergence guarantees, but typically requires unrealistic infinite-width assumptions (Jacot et al., 2018; Novak et al., 2022).
Finite-Width Extensions: Recent work has sought to extend NTK insights to finite-width settings. (Allen-Zhu et al., 2019) showed that sufficiently wide networks can maintain NTK stability, while (Arora et al., 2019) characterized when finite networks approximate the infinite-width behavior (Arora et al., 2019; Allen-Zhu et al., 2019; Frei and Gu, 2022). However, these results often require exponentially large widths to provide meaningful guarantees.
@@ 2.3 Landscape Geometry and Optimization Dynamics
Understanding the geometric structure of neural network loss landscapes has been crucial for explaining optimization success. (Choromanska et al., 2015) showed that poor local minima are exponentially rare using spin-glass models, while (Kawaguchi, 2016) proved that all local minima are global for deep linear networks.
Local Geometric Analysis: Recent work has developed general mathematical frameworks for loss landscapes in over-parameterized systems, providing modern views of efficient optimization. The LQCR framework (Aich et al., 2025) contributes to this line by providing concrete geometric conditions that ensure both convergence and fast rates.
@@
3 Notation
In this section, we introduce the notation used throughout our analysis. We consider an optimization problem for deep neural networks with the following symbols: $f_{\theta}$, $\theta$, $p$, $D$, $m$, $x$, $d$, $y$, $n$, $\mathcal{L}$, $\ell$, $\mathcal{R}$, $\eta$, $t$, $L$, $L_{\Theta}$, $\alpha$, $\gamma$, $\mu$, $\mathcal{L}^{*}$, $\mathcal{L}_{\mathcal{R}}^{*}$, $J_{\theta}$, $\boldsymbol{\theta}_{\theta}$, and $\lambda_{\min}$.
Let $f_{\theta}$ be a deep neural network with $D$ layers, width $m$, and parameters $\theta\in\mathbb{R}^{p}$. For a dataset with $n$ samples of input-label pairs $(x,y)$, where $x\in\mathbb{R}^{d}$, we aim to minimize the empirical loss function $\mathcal{L}(\theta)$, which is the average of a per-sample loss $\ell$. The gradient descent algorithm proceeds over iterations $t$ with a learning rate $\eta$. Our analysis takes place within a local region near initialization, denoted by $\mathcal{R}\subset\mathbb{R}^{p}$. Our theory connects several key concepts. The properties of a Locally Quasi-Convex Region (LQCR) are defined by a curvature parameter $\alpha$ and a descent parameter $\gamma$. The Polyak-Lojasiewicz (PL) condition is defined by a constant $\mu$, the global minimum loss value $\mathcal{L}^{*}$, and the regional minimum loss $\mathcal{L}_{\mathcal{R}}^{*}$. The loss function $\mathcal{L}$ is assumed to be $L$-smooth. Finally, our link to deep learning theory relies on the network’s Jacobian $J_{\theta}$, the resulting Neural Tangent Kernel (NTK) matrix $\boldsymbol{\theta}_{\theta}$ (which is assumed to be $L_{\Theta}$-Lipschitz), and its smallest eigenvalue, which we assume is bounded below by $\lambda_{\min}>0$ within the region $\mathcal{R}$.
4 From Local Geometry to the Polyak-Lojasiewicz Condition
The PL condition, first introduced by Polyak (Polyak, 1963), is a powerful tool for proving linear convergence on non-convex functions. However, the global PL condition, defined as
@@ 4 From Local Geometry to the Polyak-Lojasiewicz Condition
Definition 4.1 (Locally Polyak-Lojasiewicz (LPL) Region).
@@
[…]
@@
The existence of such an LPLR is not arbitrary; it is a direct consequence of the underlying geometry of the region $\mathcal{R}$. Specifically, this analysis builds upon the concept of a Locally Quasi-Convex Region (LQCR) introduced in (Aich et al., 2025). An LQCR is a region with bounded negative curvature and a guaranteed descent direction, properties that ensure iterates make steady progress.
Definition 4.2 (Locally Quasi-Convex Region (Aich et al., 2025)).
@@
  1. 1.
[…]
  2. 2.
[…]
The critical link between the geometry of an LQCR and the functional property of an LPLR is the Neural Tangent Kernel (NTK) (Jacot et al., 2018). The empirical NTK, $\Theta_{\theta}$, is defined from the network Jacobian $J_{\theta}(x)$ as $\Theta_{\theta}(x,x^{\prime})=J_{\theta}(x)J_{\theta}(x^{\prime})^{\top}$. For our analysis to hold, we assume that this kernel remains well-behaved within the LQCR. This local stability is expected in many modern deep learning scenarios, including sufficiently overparameterized or wide networks, and when using standard initialization schemes and constrained learning rates (Arora et al., 2019; Allen-Zhu et al., 2019; Aich et al., 2025).
@@
[…]
@@
Remark 4.3 (On the descent condition orientation).
In Definition 4.2, the descent condition is stated using the displacement vector $(\phi-\theta)$. This choice is without loss of generality: equivalently, one could write $(\theta-\phi)$ with the inequality sign flipped. The essential requirement is that, whenever $\mathcal{L}(\phi)<\mathcal{L}(\theta)$, the gradient at $\theta$ has a positive projection along the displacement toward $\phi$. This convention ensures the geometry is consistent and avoids confusion about sign. For consistency with this convention, we restate the descent condition from prior LQCR work (Aich et al., 2025); the equivalent form depends on whether the displacement is taken as $(\phi-\theta)$ or $(\theta-\phi)$.
5 Main Theoretical Results